OpenStudy (anonymous):

please show me how to distribute this problem to find the answer 3y-2(y+5)-6/2 btw the / sign means divided by

OpenStudy (anonymous):

u distribute the -2 to y and 5 so it is 3y-2y-10-3

OpenStudy (anonymous):

[3y-2(y+5)-6]/2 = [3y -2y+10 -6]/2 = [y+4]/2 =(y/2) +2